# Supporting

### v1

@@ Abstract
Quantile optimal treatment regimes (OTRs) aim to assign treatments that maximize a specified quantile of patients’ outcomes. Compared to treatment regimes that target the mean outcomes, quantile OTRs offer fairer regimes when a lower quantile is selected, as it focuses on improving outcomes for individuals who would otherwise experience relatively poor results. In this paper, we propose a novel method for estimating quantile OTRs by reformulating the problem as a sequential classification task. This reformulation enables us to leverage the powerful machine learning technique to enhance computational efficiency and handle complex decision boundaries. We also investigate the estimation of quantile OTRs when outcomes are discrete, a setting that has received limited attention in the literature. A key challenge is that direct extensions of existing methods to discrete outcomes often lead to inconsistency and ineffectiveness issues. To overcome this, we introduce a smoothing technique that maps discrete outcomes to continuous surrogates, enabling consistent and effective estimation. We provide theoretical guarantees to support our methodology, and demonstrate its superior performance through comprehensive simulation studies and real-data analysis.
@@ 1 Introduction
Optimal treatment regimes (OTRs) aim to assign treatments based on patients’ characteristics to achieve the best possible outcomes. Previous work on estimating OTRs has primarily focused on optimizing the mean of the outcomes, commonly referred to as “mean OTRs”. Common methods include Q-learning Watkins and Dayan, (1992), A-learning Murphy, (2003); Shi et al., 2018a , value search methods Zhang et al., 2012b , and classification methods Zhao et al., (2012); Zhou et al., (2017); Zhang et al., 2012a . Despite the popularity of the mean criterion, other criteria such as the quantile may be more appropriate in many applications. For example, the empirical mean can be heavily influenced by extreme values. In such cases, the median is often preferred due to its robustness to outliers and heavy-tailed distributions. Moreover, in situations where outcomes are measured on an ordinal scale, such as cancer stages from 1 to 4, that support ranking but not arithmetic operations like averaging, the quantile criterion offers a more appropriate and interpretable measure. OTRs that optimize a specific quantile, referred to as “quantile OTRs”, can promote fairness, particularly when targeting lower quantiles (Wang et al.,, 2018; Fang et al.,, 2023). By maximizing a lower quantile, such regimes aim to improve outcomes for individuals who would otherwise experience relatively poor results.
@@
With these insights, Linn et al., (2017) studied a general framework for estimating the quantile OTRs. Their method depends on model-based interactive Q-learning and therefore highly relies on a correctly specified model for the conditional distribution of the outcome given treatment and covariates. Later, Wang et al., (2018) studied the problem via value search methods, which established a consistent estimator of the counterfactual quantile given any regime and derived optimal regimes by maximizing it. Their method can achieve the doubly robust property by leveraging both the propensity score model and the conditional quantile function of the outcome given treatment and covariates. Along this line, Jiang et al., (2017); Zhou et al., (2023) further studied quantile OTRs for survival data.
@@
Wang et al., (2018) open the door to the application of the model-free method to quantile OTRs. However, the method exhibits several challenges. First, the optimization problem involved in the method is not concave, so there is no guarantee that the global maximizer can be derived even with the state-of-art genetic algorithm Kramer, (2017). Second, the method primarily focuses on linear regimes. While linear regimes are interpretable, they may fail to capture the potentially complex and nonlinear structure of quantile OTRs in practice, thereby limiting the flexibility of the approach. Third, when the outcomes are discrete, the method encounters issues of inconsistency and ineffectiveness. Specifically, the estimated quantile OTR may not consistently recover the true quantile OTR in terms of its corresponding quantile value across a wide range of quantile levels. Moreover, the expected counterfactual outcome under the estimated quantile OTR may not be the highest among all possible quantile OTRs. These challenges are discussed in detail in Section 2.1.
@@
To address these challenges, we propose a novel method called Sequential Classification Learning (SCL), which reformulates the problem of estimating quantile OTRs as a sequence of classification tasks solvable via convex optimization, thereby circumventing the non-convexity issue. To accommodate nonlinear treatment regimes, we allow SCL to incorporate a Gaussian kernel, allowing for flexible and expressive decision boundaries. For discrete outcomes, we address inconsistency and ineffectiveness by introducing a smoothing step that maps discrete outcomes to continuous surrogates, allowing SCL to handle discrete and continuous outcomes within a unified framework. We summarize the main contributions of this work as follows.
@@
First, we propose a novel method that leverages the classification learning to estimate quantile OTRs. The classification-based method has demonstrated notable strengths in estimating mean OTRs Zhang et al., 2012a , as they can harness machine learning techniques to enhance computational efficiency, avoid issues of local maximizers, and flexibly accommodate nonlinear treatment regimes. Motivated by these advantages, we seek to extend classification learning to the quantile setting. However, a direct extension is not feasible because the quantile function is not a linear operator. To overcome this challenge, we develop a new reformulation that recasts the estimation of quantile OTRs as a sequential classification problem, guided by an adaptive binary search procedure. Leveraging this novel reformulation, we propose a doubly robust estimator for quantile OTRs by incorporating both the propensity score and the conditional survival function of the outcome given treatment and covariates. We further extend the method to dynamic treatment settings, where treatment decisions evolve over time based on patients’ intermediate responses to prior treatments.
@@
Third, we provide theoretical guarantees for our method. Our approach involves a sequence of interdependent classification tasks guided by a binary search procedure, posing significant analytical challenges due to the recursive dependence structure. To the best of our knowledge, asymptotic theory for such binary search-based procedures is largely unexplored. We address this gap by employing empirical process techniques. Our findings demonstrate that, when subjects with differing optimal treatments are well separated, the convergence rate of the quantile under the estimated regime to that under the true quantile OTR approaches $n^{-1/3}$ when employing a Gaussian kernel. Moreover, the convergence rate of our method is close to that of an oracle method, one possessing prior knowledge of the optimal quantile. En route to establishing these asymptotic results, we also derive novel theoretical insights for classification learning combined with the Gaussian kernel, thereby addressing gaps previously unfilled in Zhang et al., 2012a .
The remaining sections of this paper are organized as follows. Section 2 presents the main methodology. Theoretical properties for SCL are provided in Section 3. The finite sample performance is explored through simulations in Section 4. Section 5 illustrates the application using the ACTG175 dataset. The extension to quantile dynamic OTRs is provided in Section 6. We end with a discussion in Section 7.
@@ 2 Methodology
Assumption 1: (i) $Y=Y(0)I(A=0)+Y(1)I(A=1)$; (ii) $Y(0),Y(1)\perp\!\!\!\perp A|\boldsymbol{X}$; (iii) $10<P(A=1|\boldsymbol{X})<$ 1.
In this paper, our primary interest is the quantile OTRs (Wang et al.,, 2018) defined as:
@@
where $\pi^{*}(a|\boldsymbol{x})=P(A=a|\boldsymbol{X}=\boldsymbol{x})$ is the propensity score. We utilize the survival function to define the quantile, as our novel method, SCL, is inspired by this definition, and such a quantile definition is equivalent to the traditional definition using the cumulative distribution function. See Section LABEL:s6.2 of the supplementary material for details.
@@ 2.1 Challenges of the value search method
Suppose the observed data $\{\boldsymbol{X}_{i},Y_{i},A_{i}\},i=1,\ldots,n$ are independent and identically distributed copies of $\{\boldsymbol{X},Y,A\}$. To estimate the quantile OTRs, Wang et al., (2018) first estimate $Q_{\tau}\{Y(d)\}$ by
@@
where $\rho_{\tau}\left(u\right)=u\{\tau-I(u<0)\}$ is the check loss. Then they consider the set of linear ITRs $\mathcal{D}=\{d(\boldsymbol{x})=I(\beta_{0}+{\boldsymbol{x}}^{T}\boldsymbol{\beta}\geq 0):|\beta_{0}|=0,1\text{ and }\boldsymbol{\beta}\in\mathbb{R}^{d}\}$, and estimate quantile OTRs by maximizing $\widehat{Q}_{\tau}\{Y(d)\}$ over the set $\mathcal{D}$ based on the genetic algorithm (Kramer,, 2017).
The value search method proposed by Wang et al., (2018) lays a strong foundation for the estimation of quantile OTRs. However, it exhibits several challenges:
  1. 1.
    Non-convexity: The objective function $\widehat{Q}_{\tau}\{Y(d)\}$ is not concave with respect to the parameters $(\beta_{0},\boldsymbol{\beta}^{T})^{T}$. As a result, the solution based on the genetic algorithm may correspond to a local maximizer, potentially leading to a harmful treatment regime, and posing significant risks in real-world applications. This issue becomes more critical when $\boldsymbol{\beta}$ becomes high dimensional (Shi et al., 2018b, ). We provide a simulation in Section LABEL:s4.1 of the supplementary material. In that particular setting, Wang et al., (2018)’s method fails to find the global maximizer $61.5$% of the time.
  2. 2.
    Regime linearity: Wang et al., (2018) mainly studied the linear regime. While the linear regime is easy to interpret, in practice, there may be greater interest in more flexible, nonlinear quantile OTRs. For example, consider a scenario where the outcome model follows a simple heteroscedastic linear form, i.e., $Y=m(\boldsymbol{X})+Ag(\boldsymbol{X})+h(\boldsymbol{X},A)\epsilon$, where $\epsilon$ is a random noise and $m(\boldsymbol{X})$, $g(\boldsymbol{X})$, and $h(\boldsymbol{X},A)$ are all linear functions. In this case, the mean OTRs are linear, however, the quantile OTRs are nonlinear. A detailed derivation of this result is provided in Section LABEL:s6.4 of the supplementary material. This highlights that even simple linear outcome models can yield nonlinear quantile OTRs, emphasizing the importance of exploring nonlinear approaches for quantile OTRs.
  3. 3.
    Inconsistency: In the OTR literature, value consistency is a commonly established property (Zhao et al., 2015b, ). For continuous outcomes, Wang et al., (2018)’s method enjoys quantile value consistency; i.e., $Q_{\tau}\{Y(\hat{d})\}$ converges in probability to $\max_{d\in\mathcal{D}}Q_{\tau}\{Y(d)\}$. However, when their method is applied directly to discrete outcomes, this consistency property no longer holds. We demonstrate this through a toy simulation study in Section LABEL:s4.2 of the supplementary material. We compute the empirical mean squared error (MSE) obtained by Wang et al., (2018)’s method, defined as the average of $[Q_{\tau}\{Y(\hat{d})\}-\max_{d\in\mathcal{D}}Q_{\tau}\{Y(d)\}]^{2}$ across $200$ Monte Carlo replications. The results are presented in Figure 1, which shows that the empirical MSE of Wang et al., (2018)’s method does not diminish as the sample size increases across various quantile levels. In contrast, our proposed method exhibits a decreasing MSE when sample size increases, providing evidence of consistency. Notably, the MSE of our method can reach zero due to the discrete nature of the outcome. Full simulation details and an extended explanation of these results are provided in Section LABEL:s4.2 of the supplementary material. The inconsistency issue stems from the fact that the quantile estimator (2) is not a consistent estimator of the quantile $Q_{\tau}\{Y(d)\}$ when the outcome is discrete Machado and Silva, (2005). As a consequence, maximizing (2) does not necessarily yield a consistent estimator of the quantile OTRs.
@@
    Ineffectiveness: Define $\mathcal{D}^{*}=\{\tilde{d}\in\mathcal{D}:Q_{\tau}\{Y(\tilde{d})\}=\max_{d\in\mathcal{D}}Q_{\tau}\{Y({d})\}\}$ as the set of quantile OTRs. Since $\mathcal{D}^{*}$ typically comprises multiple regimes in the case of discrete outcomes, it may be desirable to identify the one that best satisfies an additional criterion. In this paper, we introduce a natural criterion that compares their expected counterfactual outcomes, $E\{Y(d^{*})\}$ for each $d^{*}\in\mathcal{D}^{*}$, and selects the regime with the highest expected counterfactual outcomes. We refer to such a regime as an effective quantile OTR. However, the quantile OTR obtained by solving (1) (Wang et al.,, 2018) is not guaranteed to be effective under this criterion. Therefore, estimating quantile OTRs via maximizing the empirical version of $Q_{\tau}\{Y(d)\}$, i.e., $\hat{d}=\operatorname*{arg\,max}_{d\in\mathcal{D}}\widehat{Q}_{\tau}\{Y(d)\}$, also suffers from the ineffectiveness issue. We have provided details in Section LABEL:s3.4 of the supplementary material.
2.2 Sequential classification learning for quantile OTRs with continuous outcome
To address these challenges, we propose a novel approach inspired by classification learning. Classification learning methods have demonstrated notable strengths in estimating mean OTRs Zhang et al., 2012a ; Zhao et al., (2012); for example, they can leverage machine learning techniques to accelerate computation, avoid issues of local maximizers, and accommodate flexible, nonlinear regimes. However, they cannot be directly extended to quantile OTRs because the quantile function is not a linear operator. To overcome this limitation, we introduce a new reformulation that transforms the estimation of quantile OTRs into a sequential classification learning problem. We refer to our method as Sequential Classification Learning (SCL).

### v2

# Successive classification learning for estimating quantile optimal treatment regimes

@@ Abstract
Quantile optimal treatment regimes (OTRs) aim to assign treatments that maximize a specified quantile of patients’ outcomes. Compared to treatment regimes that target the mean outcomes, quantile OTRs offer fairer regimes when a lower quantile is selected, as it improves outcomes for vulnerable patients. In this paper, we propose a novel method for estimating quantile OTRs by reformulating the problem as a successive classification task, solvable via training a sequence of classifiers, each successive classifier built on the output of its predecessors. This reformulation enables us to leverage the powerful machine learning technique to enhance computational efficiency and handle complex decision boundaries. We also investigate the estimation of quantile OTRs when outcomes are discrete, a setting that has received limited attention in the literature. A key challenge is that direct extensions of existing methods to discrete outcomes often lead to inconsistency and ineffectiveness issues. To overcome this, we introduce a smoothing technique that maps discrete outcomes to continuous surrogates, enabling consistent and effective estimation. We provide theoretical guarantees to support our methodology, and demonstrate its superior performance through comprehensive simulation studies and real-data analysis.
@@ 1 Introduction
Optimal treatment regimes (OTRs) aim to assign treatments based on patients’ characteristics to achieve the best possible outcomes. Previous work on estimating OTRs has primarily focused on optimizing the mean of the outcomes, commonly referred to as “mean OTRs.” Common methods include Q-learning watkins1992qlearning, A-learning murphy2003optimal; shi2018highdimensional, value search methods zhang2012a, and classification methods zhao2012estimating; zhou2017residual; zhang2012estimating. Despite the popularity of the mean criterion, other criteria such as the quantile may be more appropriate in many applications. For example, the empirical mean can be heavily influenced by extreme values. In such cases, the median is often preferred due to its robustness to outliers and heavy-tailed distributions. Moreover, in situations where outcomes are measured on an ordinal scale, such as cancer stages from 1 to 4, that support ranking but not arithmetic operations like averaging, the quantile criterion offers a more appropriate and interpretable measure. OTRs that optimize a specific quantile, referred to as “quantile OTRs”, can promote fairness, particularly when targeting lower quantiles (wang2018quantileoptimal; fang2023fairness). Specifically, patients whose outcomes fall in the lower quantiles of the distribution are considered more vulnerable, as they experience relatively poorer results compared to the broader population. By explicitly targeting improvements in these lower quantiles, quantile OTRs aim to enhance outcomes for vulnerable patients and thereby promote fairness.
@@
With these insights, linn2017interactive studied a general framework for estimating the quantile OTRs. Their method depends on model-based interactive Q-learning and therefore highly relies on a correctly specified model for the conditional distribution of the outcome given treatment and covariates. Later, wang2018quantileoptimal studied the problem via value search methods, which established a consistent estimator of the counterfactual quantile given any regime and derived optimal regimes by maximizing it. Their method can achieve the doubly robust property by leveraging both the propensity score model and the conditional quantile function of the outcome given treatment and covariates. Along this line, jiangDoublyRobustEstimation2017; zhouTransformationInvariantLearningOptimal2022 further studied quantile OTRs for survival data.
@@
wang2018quantileoptimal open the door to the application of the model-free method to quantile OTRs. However, the method exhibits several challenges. First, the optimization problem involved in the method is not concave, so there is no guarantee that the global maximizer can be derived even with the state-of-art genetic algorithm kramer2017genetic. Second, the method primarily focuses on linear regimes. While linear regimes are interpretable, they may fail to capture the potentially complex and nonlinear structure of quantile OTRs in practice, thereby limiting the flexibility of the approach. Third, when the outcomes are discrete, the method encounters issues of inconsistency and ineffectiveness. Specifically, the estimated quantile OTR may not consistently recover the true quantile OTR in terms of its corresponding quantile value across a wide range of quantile levels. Moreover, the expected counterfactual outcome under the estimated quantile OTR may not be the highest among all possible quantile OTRs. These challenges are discussed in detail in Section 2.1.
@@
To address these challenges, we propose a novel method called Successive Classification Learning (SCL), which reformulates the problem of estimating quantile OTRs as a sequence of classification tasks solvable via convex optimization, thereby circumventing the non-convexity issue. To accommodate nonlinear treatment regimes, we allow SCL to incorporate a Gaussian kernel, allowing for flexible and expressive decision boundaries. For discrete outcomes, we address inconsistency and ineffectiveness by introducing a smoothing step that maps discrete outcomes to continuous surrogates, allowing SCL to handle discrete and continuous outcomes within a unified framework. We summarize the main contributions of this work as follows.
@@
First, we propose a novel method that leverages the classification learning to estimate quantile OTRs. The classification-based method has demonstrated notable strengths in estimating mean OTRs zhang2012estimating, as they can harness machine learning techniques to enhance computational efficiency, avoid issues of local maximizers, and flexibly accommodate nonlinear treatment regimes. Motivated by these advantages, we seek to extend classification learning to the quantile setting. However, a direct extension is not feasible because the quantile function is not a linear operator. To address this challenge, we recast the estimation of quantile OTRs as a successive classification problem, solvable by training a sequence of classification models, each one building upon the output of its predecessors and guided by a binary search procedure to refine the quantile OTR estimates. Leveraging this novel reformulation, we propose a doubly robust estimator for quantile OTRs by incorporating both the propensity score and the conditional survival function of the outcome given treatment and covariates. To enrich the applicability of our framework, we further extend our method to dynamic treatment settings, in which treatment decisions adapt over time according to patients’ intermediate responses to prior treatments, and to survival data, where outcomes of interest are subject to right censoring. These extensions are presented in Section S4 of the supplementary material.
@@
Third, we provide theoretical guarantees for our method. Our approach involves a sequence of interdependent classification tasks guided by a binary search procedure, posing significant analytical challenges due to the recursive dependence structure. To the best of our knowledge, asymptotic theory for such binary search-based procedures is largely unexplored. We address this gap by employing empirical process techniques. Our findings demonstrate that, when subjects with differing optimal treatments are well separated, the convergence rate of the quantile under the estimated regime to that under the true quantile OTR approaches $n^{-1/3}$ when employing a Gaussian kernel. Moreover, the convergence rate of our method is close to that of an oracle method, one possessing prior knowledge of the optimal quantile. En route to establishing these asymptotic results, we also derive novel theoretical insights for classification learning combined with the Gaussian kernel, thereby addressing gaps previously unfilled in zhang2012estimating.
The remaining sections of this paper are organized as follows. Section 2 presents the main methodology. Theoretical properties for SCL are provided in Section 3. The finite sample performance is explored through simulations in Section 4. Section 5 illustrates the application using the ACTG175 dataset. We end with a discussion in Section 6.
@@ 2 Methodology
Assumption 1: (i) $Y=Y(0)I(A=0)+Y(1)I(A=1)$; (ii) $Y(0),Y(1)\perp\!\!\!\perp A|\boldsymbol{X}$; (iii) $\epsilon<P(A=1|\boldsymbol{X})<1-\epsilon$ for some $\epsilon>0$.
Assumption 1 is a fundamental assumption in causal inference to ensure identification hernán2020causal. Assumption 1(i) is referred to as the stable unit treatment value assumption (SUTVA), which requires each patient’s outcome depends solely on patient’s own treatment. Assumption 1(ii) is referred to as the ignorability assumption, which guarantees that, conditional on $\boldsymbol{X}$, treatment assignment $A$ is independent of the potential outcomes. Assumption 1(iii) is referred to as the positivity assumption, which means that every subgroup defined by covariates has a positive probability of receiving each treatment option.
In this paper, our primary interest is the quantile OTRs (wang2018quantileoptimal) defined as:
@@
where $\pi^{*}(a|\boldsymbol{x})=P(A=a|\boldsymbol{X}=\boldsymbol{x})$ is the propensity score. We utilize the survival function to define the quantile, as our novel method, SCL, is inspired by this definition, and such a quantile definition is equivalent to the traditional definition using the cumulative distribution function. See Section S5.1 of the supplementary material for details.
@@ 2.1 Challenges of the value search method
Suppose the observed data $\{\boldsymbol{X}_{i},Y_{i},A_{i}\},i=1,\ldots,n$ are independent and identically distributed copies of $\{\boldsymbol{X},Y,A\}$. To estimate the quantile OTRs, wang2018quantileoptimal first estimate $Q_{\tau}\{Y(d)\}$ by
$\widehat{Q}_{\tau}\{Y(d)\}=\underset{q}{\operatorname*{arg\,min}}\frac{1}{n}\sum_{i=1}^{n}\frac{I\{d(\boldsymbol{X}_{i})=A_{i}\}}{\pi^{*}(A_{i}|\boldsymbol{X}_{i})}\rho_{\tau}\left(Y_{i}-q\right),$ (2)
where $\rho_{\tau}\left(u\right)=u\{\tau-I(u<0)\}$ is the check loss. Then they consider the set of linear ITRs $\mathcal{D}=\{d(\boldsymbol{x})=I(\beta_{0}+{\boldsymbol{x}}^{T}\boldsymbol{\beta}>0):|\beta_{0}|=0,1\text{ and }\boldsymbol{\beta}\in\mathbb{R}^{d}\}$, and estimate quantile OTRs by maximizing $\widehat{Q}_{\tau}\{Y(d)\}$ over the set $\mathcal{D}$ based on the genetic algorithm (kramer2017genetic).
The value search method proposed by wang2018quantileoptimal lays a strong foundation for the estimation of quantile OTRs. However, it exhibits several challenges:
  1. 1.
    Non-convexity: The objective function $\widehat{Q}_{\tau}\{Y(d)\}$ is not concave with respect to the parameters $(\beta_{0},\boldsymbol{\beta}^{T})^{T}$. As a result, the solution based on the genetic algorithm may correspond to a local maximizer, potentially leading to a harmful treatment regime, and posing significant risks in real-world applications. This issue becomes more critical when $\boldsymbol{\beta}$ becomes high dimensional (shi2018forward). We provide a simulation in Section S1.1 of the supplementary material. In that particular setting, wang2018quantileoptimal’s method fails to find the global maximizer 61.561.5% of the time.
  2. 2.
    Regime linearity: wang2018quantileoptimal mainly studied the linear regime. While the linear regime is easy to interpret, in practice, there may be greater interest in more flexible, nonlinear quantile OTRs. For example, consider a scenario where the outcome model follows a simple heteroscedastic linear form, i.e., $Y=m(\boldsymbol{X})+Ag(\boldsymbol{X})+h(\boldsymbol{X},A)\epsilon$, where $\epsilon$ is a random noise and $m(\boldsymbol{X})$, $g(\boldsymbol{X})$, and $h(\boldsymbol{X},A)$ are all linear functions. In this case, the mean OTRs are linear, however, the quantile OTRs are nonlinear. A detailed derivation of this result is provided in Section S1.2 of the supplementary material. This highlights that even simple linear outcome models can yield nonlinear quantile OTRs, emphasizing the importance of exploring nonlinear approaches for quantile OTRs.
  3. 3.
@@
    $Q_{\tau}\{Y(d)\}$. However, when their method is applied directly to discrete outcomes, this consistency property no longer holds. We demonstrate this through a toy simulation study in Section S1.3 of the supplementary material. We compute the empirical mean squared error (MSE) obtained by wang2018quantileoptimal’s method, defined as the average of $Q_{\tau}\{Y(\hat{d})\}-\max_{d\in\mathcal{D}}Q_{\tau}\{Y(d)\}]^{2}$ across $200$ Monte Carlo replications. The results are presented in Figure 1, which shows that the empirical MSE of wang2018quantileoptimal’s method does not diminish as the sample size increases across various quantile levels. In contrast, our proposed method exhibits a decreasing MSE when sample size increases, providing evidence of consistency. Notably, the MSE of our method can reach zero due to the discrete nature of the outcome. Full simulation details and an extended explanation of these results are provided in Section S1.3 of the supplementary material. The inconsistency issue stems from the fact that the quantile estimator (2) is not a consistent estimator of the quantile $Q_{\tau}\{Y(d)\}$ when the outcome is discrete machado2005quantiles. As a consequence, maximizing (2) does not necessarily yield a consistent estimator of the quantile OTRs.
@@
    Ineffectiveness: Define $\mathcal{D}^{*}=\{\tilde{d}\in\mathcal{D}:Q_{\tau}\{Y(\tilde{d})\}=\max_{d\in\mathcal{D}}Q_{\tau}\{Y({d})\}\}$ as the set of quantile OTRs. Since $\mathcal{D}^{*}$ typically comprises multiple regimes in the case of discrete outcomes, it may be desirable to identify the one that best satisfies an additional criterion. In this paper, we introduce a natural criterion that compares their expected counterfactual outcomes, $E\{Y(d^{*})\}$ for each $d^{*}\in\mathcal{D}^{*}$, and selects the regime with the highest expected counterfactual outcomes. We refer to such a regime as an effective quantile OTR. However, the quantile OTR obtained by solving (1) (wang2018quantileoptimal) is not guaranteed to be effective under this criterion. Therefore, estimating quantile OTRs via maximizing the empirical version of $Q_{\tau}\{Y(d)\}$, i.e., $\hat{d}=\operatorname*{arg\,max}_{d\in\mathcal{D}}\widehat{Q}_{\tau}\{Y(d)\}$, also suffers from the ineffectiveness issue. We have provided details in Section S1.4 of the supplementary material.
2.2 Successive classification learning for continuous outcomes
To address these challenges, we propose a novel approach inspired by classification learning. Classification learning methods have demonstrated notable strengths in estimating mean OTRs zhang2012estimating; zhao2012estimating; for example, they can leverage machine learning techniques to accelerate computation, avoid issues of local maximizers, and accommodate flexible, nonlinear regimes. However, they cannot be directly extended to quantile OTRs because the quantile function is not a linear operator. To overcome this limitation, we recast the estimation of quantile OTRs as a successive classification problem, solvable through training a sequence of classifiers, each building on the outputs of its predecessors. We refer to our method as Successive Classification Learning (SCL).